# Repairs in a Block World: A New Benchmark for Handling User Corrections with Multi-Modal Language Models

### v1

@@ Abstract
In dialogue, the addressee may initially misunderstand the speaker and respond erroneously, often prompting the speaker to correct the misunderstanding in the next turn with a Third Position Repair (TPR). The ability to process and respond appropriately to such repair sequences is thus crucial in conversational AI systems. In this paper, we first collect, analyse, and publicly release BlockWorld-Repairs: a dataset of multi-modal TPR sequences in an instruction-following manipulation task that is, by design, rife with referential ambiguity. We employ this dataset to evaluate several state-of-the-art Vision and Language Models (VLM) across multiple settings, focusing on their capability to process and accurately respond to TPRs and thus recover from miscommunication. We find that, compared to humans, all models significantly underperform in this task. We then show that VLMs can benefit from specialised losses targeting relevant tokens during fine-tuning, achieving better performance and generisability. Our results suggest that these models are not yet ready to be deployed in multi-modal collaborative settings where repairs are common, and highlight the need to design training regimes and objectives that facilitate learning from interaction.
@@ 1 Introduction
Figure 1: Example dialogue from BlockWorld-Repairs: after predicting an incorrect response, VLMs have to be able to interpret the repair in order to produce the correct bounding box prediction, a critical skill for human-robot collaboration tasks.
@@
Therefore, the ability to interpret and generate effective repair sequences is crucial to robust and faithful Conversational AI technology. This need is especially acute in settings where more fine-grained levels of understanding are required, for example in embodied human-machine collaboration. Indeed the frequency of miscommunication in human dialogue is known to vary with both the overarching task as well as the medium of communication Colman and Healey (2011).
@@
In this paper, we focus on a particular class of corrective feedback, so called, Third Position Repairs (henceforth TPR; Schegloff, 1992) in multi-modal settings. These occur when the addressee initially misunderstands the speaker (see Figure 1 at T1, the trouble source turn), responds based on this misunderstanding (at T2), which in turn reveals the misunderstanding to the addressee who then goes on to correct the misunderstanding (at T3). The addressee then provides a new response (at T4) to the original request (at T1) based on the correction. TPRs have been largely neglected in the NLP community, likely due to underestimation of the importance of repair phenomena in dialogue, and thereby lack of appropriate data and frameworks for model training and evaluation. Repair is of particular interest in (multi-modal) language modelling since it almost invariably involves highly focused / specific revisions to the representations already built by the model (e.g. revisions to the referents of referring expressions, like in our case here), based on the context-dependent structure of the corrective feedback. Repairs can thus be used as probes into how models represent meaning and the structure of the overarching task (see Chiyah-Garcia et al. (2023); Madureira and Schlangen (2023); Benotti and Blackburn (2021) for different implementations of this idea). And as we show here, this is also consequential for how language models should be trained to process TPRs including the specific objective functions involved. Our contributions are thus as follows: 1) we collect and release BlockWorld-Repairs (BW-R), a dataset of collaborative dialogues in a table-top manipulation task that, by design, has a high potential for referential ambiguity and is focused on complex multi-modal task instructions as well as TPRs; 2) we establish a human baseline for the whole task through an in-person human study where participants try to solve the task using these dialogues; 3) We show that state-of-the-art Vision Language Models (VLM) can learn to process TPR sequences through a specialized training regime, with some opportunity for generalisation; 4) Considering the substantial performance gap between the best models and humans, we present an in-depth error analysis comparing models and humans.
@@ 2 Background
Considerable attention has been paid to computational models for the interpretation and generation of self-repair (see Hough and Schlangen, 2015; Hough, 2015; Shalyminov et al., 2017; Eshghi and Ashrafzadeh, 2023; Buß and Schlangen, 2011; Hough and Purver, 2012, among others): a class of repairs whereby the speaker corrects themselves on the fly within the same conversational turn (e.g. “User: I want to go to London uhm sorry Paris”). Similarly, the crucial role of generating and responding to CRs (e.g. “Pardon/what/who?”) in conversational models has long been recognised (see San-Segundo et al., 2001; Purver, 2004; Purver and Ginzburg, 2004; Rieser and Moore, 2005; Rodríguez and Schlangen, 2004; Rieser and Lemon, 2006, among others), but existing systems either remain limited (e.g. Curry et al. (2018)) or do not support this at all – see Purver et al. (2018) for an overview. Recent work tries to identify when to pose a CR Madureira and Schlangen (2023); Zhu et al. (2021); Shi et al. (2022), but few evaluate the ability of models to process their responses Gervits et al. (2021); Aliannejadi et al. (2021). Chiyah-Garcia et al. (2023) show that the ability of Vision and Language Models (VLM) to effectively process clarificational exchanges depends on the level of granularity of the model’s cross-modal representation, and therefore also the specific objectives used to train the models.
@@ 3 The BlockWorld-Repair Dataset
Prior works have struggled to handle the most ambiguous setting with blank blocks Tan and Bansal (2018); Mehta and Goldwasser (2019); Dan et al. (2021). However, here we argue that single-turn instructions are not sufficient to solve highly ambiguous environments or real-world situations. Similar to how humans misinterpret and subsequently repair referential ambiguities with TPRs, VLMs should also be able to handle multi-modal TPRs in dialogues.
@@
In contrast to more recent photo-realistic simulated environments (e.g., AI2Thor Kolve et al. (2017)), we build BW-R using the Unity engine to simulate a realistic tabletop manipulation task that gives us the ability to specifically control for the conditions that require collaboration between agents to correctly complete the task. This is also similar to previous works Dobnik et al. (2020); Jiang et al. (2023), e.g., VIMABench sacrifices photo-realistic vision to extensively assess the systematic generalisation ability of VLMs for tabletop manipulation tasks.
@@ 3.1 Dataset Collection
As a result of this data collection, we obtained 795 dialogues after filtering out low-quality and problematic ones, which include 795 initial instructions, 629 source block TPRs and 635 final target position TPRs, resulting in 2059 total entries. We provide further details about the AMT setup, task and dataset statistics in Section A.1.
@@ 4.1.1 Dataset
We use the BW-R data (70/30 train/test), which contains two subsets: single-instructions (1 turn) and TPRs (3 turns: instruction, initial system response, repair). We ablate the data during our experiments to analyse fine-grained information about the models’ capabilities. Refer to Appendix A for further dataset details.
@@ 4.1.2 Tasks
Previous works have formulated these tasks to aid the prediction, such as finding a reference block and an offset distance from it Bisk et al. (2016) or quadrants in a board Mehta and Goldwasser (2019). For our experiments, and following common visual grounding objectives in VLMs Wang et al. (2022); Liu et al. (2024); Chen et al. (2023), we prompt models to generate the coordinates of the bounding box for the block or location ($[x_{min},y_{min},x_{max},y_{max}]$) in the range of 0-1, which we then scale to the image size. Predictions not matching this format during parsing are marked as out of distribution and removed. We use specific prompts for each model to match their training prompting strategy (see Appendix B).
@@ Results
Overall, the models struggle with handling TPRs, especially in zero-shot settings. Intuitively, repairs should facilitate the task by introducing new information that refines the candidate pool. GPT-4o is unique in its ability to process these repairs effectively and improve its performance with instructions. We see that models can learn to process TPRs with some training, and the following section explores ways to further distil this capability.
@@ 5.2 Results
Table 2 shows the fine-tuned VLMs on the full data across different masking criteria [^4] All model results are provided in Appendix D.. We first observe that applying a mask on all tokens but completions has large benefits for both VLMs for repairs in source prediction, with strong improvements particularly when fine-tuning with repairs-only or the full dataset. This loss allows models to learn from TPRs and transfer some task knowledge from instructions to repairs when using the full data. In other settings, this loss has small negative effects on the VLM’s performance, especially when models are trained and tested on the same data. This is somewhat expected as calculating the loss for the complete input has benefits, such as using many more tokens during training, learning the input format and quickly adapting to the task. However, these models do not seem to learn the task well and thus do not generalise.
@@
Furthermore, we see that models fine-tuned on instructions alone, with or without masking, acquire some task capabilities, but do not generalise well to corrections, and vice-versa. When trained with the full data and completion-only loss, VLMs learn what is important regardless of the conversation, and reach their best results, learning what is important particularly for repairs.
@@ 6.1 Human Comparison
To compare models with humans on this task, we use the data samples that we collected from our in-person human study, in which participants saw the same data that models are evaluated on. The results in Table 3 show clear differences, with models struggling to reach human performance. We can see that, despite Idefics2 and LLaVA having lower source accuracy for instructions, they both beat GPT-4o on distance for source. In this case, accuracy fails as a metric for prediction quality, since lower distances better reflect true proximity. We also see impressive results for LLaVA processing TPRs after our training for source prediction. Regarding target position, this task is considerably more challenging for humans and models, reflected in worse scores overall. LLaVA and Idefics2 struggle to process repairs related to target locations, in contrast to GPT-4o which outperforms the human participants.
@@ 6.2 Impact of Task Difficulty
To better pinpoint the strengths and weaknesses of the different models, we categorise the BW-R data into difficulty levels according to human performance. We use lower human performance as an indirect proxy for the complexity of the example, and define three levels as follows: 1) Easy: both human annotators for the same entry correctly predicted the block (100% source accuracy) or their target position was within 1 block unit distance away from the true location; 2) Medium: at least one out of the two annotators correctly predicted the block (50% source accuracy) or their target position was below the human’s mean distance away from the location (between 1 and 3.22 units away); 3) Hard: neither of the annotators correctly predicted the block (0% source accuracy) or their target position was further than the human’s mean distance (above 3.22 units).
@@
The results of evaluating VLMs on these difficulty subsets (Table 4) indicate that models and humans do not struggle on the same entries. Models show a more uniform distribution of performance, with improvements over humans in the most difficult subset but significantly underperforming with easier references. This behaviour seems to follow the trend with the average word number in the subsets [^5] See Appendix C, suggesting that easy examples (most words) may employ longer, more complex terms to describe references, whereas hard examples (fewest words) may have easier terms for models, but are too underspecified for humans.
@@ 6.3 Qualitative Analysis
We further analyse 50 BW-R dialogues to understand the proficiency differences of VLMs with that of human participants (see Figure 3 for some examples). We find that humans and GPT-4o tend to have similar predictions when processing candidate responses and their repairs. Models are usually able to process simple spatial repairs (i.e., left, right, above), however, they particularly struggle with more abstract concepts (i.e., rows, columns). In that regard, both Idefics2 and LLaVA exhibit comprehension of the simpler referring expressions, but have a challenging time with longer or more complex sequences. Idefics2 also seems prone to over-correct (right pushes predictions to the absolute right). We also notice that VLMs often predict bounding boxes that do not align with blocks in the source prediction task, which is unexpected given their pre-training (and fine-tuning) with referring expression objectives. For the target prediction task, VLMs significantly struggle and do not process repairs. GPT-4o is the only exception, which shows a more fine-grained understanding of the task yet falls behind on more difficult dialogues.
@@ 7 Conclusion
This paper explores VLM’s capabilities of handling repairs with a new dataset, BlockWorld-Repairs. We collect collaborative dialogues in a highly ambiguous tabletop manipulation environment that requires complex multi-modal task instructions as well as repairs. In these scenarios, referential ambiguities are common and thus being able to process TPRs becomes essential to complete the task. We validate these dialogues through an in-person human study where participants attempt to solve the task, providing a baseline for model performance.
@@
We finish with an error analysis comparing these models and human participants, finding stark differences in how they process repairs. Surprisingly, models struggle with the dialogues that human participants found easiest, highlighting a large gap in their ability to process TPRs. We believe that future research should focus on two major improvements for VLMs: 1) design training objectives that facilitate learning from interaction including TPRs; 2) equip models with more object-centric visual representations that facilitate visual grounding tasks which are of paramount importance for situated collaborative tasks.
@@ 8 Limitations
Another limitation of our work is that we do not attempt to model the entire spectrum of a robotic manipulation task (e.g., Team et al. (2024)). Instead, we follow previous work in Embodied AI (e.g., ALFRED Shridhar et al. (2020), Simbot Arena Gao et al. (2024), inter alia) which casts object manipulation as API call generation of an action label combined with a bounding box around the object to be manipulated. Although we don’t generate actions, the prediction tasks that we explore in BW-R involve generating the coordinates of a bounding box that can be used for picking up a specific block or putting it down onto a specific box on the board.

### v2

@@ Abstract
In dialogue, the addressee may initially misunderstand the speaker and respond erroneously, often prompting the speaker to correct the misunderstanding in the next turn with a Third Position Repair (TPR). The ability to process and respond appropriately to such repair sequences is thus crucial in conversational AI systems. In this paper, we first collect, analyse, and publicly release BlockWorld-Repairs: a dataset of multi-modal TPR sequences in an instruction-following manipulation task that is, by design, rife with referential ambiguity. We employ this dataset to evaluate several state-of-the-art Vision and Language Models (VLM) across multiple settings, focusing on their capability to process and accurately respond to TPRs and thus recover from miscommunication. We find that, compared to humans, all models significantly underperform in this task. We then show that VLMs can benefit from specialised losses targeting relevant tokens during fine-tuning, achieving better performance and generalising better to new scenarios. Our results suggest that these models are not yet ready to be deployed in multi-modal collaborative settings where repairs are common, and highlight the need to design training regimes and objectives that facilitate learning from interaction. Our code and data are available at www.github.com/JChiyah/blockworld-repairs
@@ 1 Introduction
Figure 1: Example dialogue from BlockWorld-Repairs: after predicting an incorrect response, VLMs must accurately interpret the repair to produce the correct bounding box prediction, a critical skill for human-robot collaboration tasks.
@@
Therefore, the ability to interpret and generate effective repair sequences is crucial to robust and faithful Conversational AI technology. This need is especially acute in settings where more fine-grained levels of understanding are required, for example in embodied human-machine collaboration. Indeed the frequency of miscommunication in human dialogue is known to vary with both the overarching task and the medium of communication Colman and Healey (2011).
@@
In this paper, we focus on a particular class of corrective feedback, so called, Third Position Repairs (henceforth TPR; Schegloff, 1992) in multi-modal settings. These occur when the addressee initially misunderstands the speaker (see Figure 1 at T1, the trouble source turn), responds based on this misunderstanding (at T2), which in turn reveals the misunderstanding to the addressee who then goes on to correct the misunderstanding (at T3). The addressee then provides a new response (at T4) to the original request (at T1) based on the correction. TPRs have been largely neglected in the NLP community, likely due to underestimation of the importance of repair phenomena in dialogue, and thereby lack of appropriate data and frameworks for model training and evaluation. Repair is of particular interest in (multi-modal) language modelling since it almost invariably involves highly focused / specific revisions to the representations already built by the model (e.g. revisions to the referents of referring expressions, like in our case here), based on the context-dependent structure of the corrective feedback. Repairs can thus be used as probes into how models represent meaning and the structure of the overarching task (see Chiyah-Garcia et al. (2023); Madureira and Schlangen (2023); Benotti and Blackburn (2021) for different implementations of this idea). And as we show here, this is also consequential for how language models should be trained to process TPRs including the specific objective functions involved.
Our contributions are thus as follows: 1) we collect and release BlockWorld-Repairs (BW-R), a dataset of collaborative dialogues in a tabletop manipulation task that, by design, has a high potential for referential ambiguity and is focused on complex multi-modal task instructions as well as TPRs; 2) we establish a human baseline for the whole task through an in-person human study where participants try to solve the task using these dialogues; 3) we show that state-of-the-art Vision Language Models (VLM) can learn to process TPR sequences through a specialised training regime, with some opportunity for generalisation; and 4) considering the substantial performance gap between the best models and humans, we present an in-depth error analysis comparing the two, revealing that models struggle with references that humans find easy.
@@ 2 Background
Considerable attention has been paid to computational models for the interpretation and generation of self-repair (see Hough and Schlangen, 2015; Hough, 2015; Shalyminov et al., 2017; Eshghi and Ashrafzadeh, 2023; Buß and Schlangen, 2011; Hough and Purver, 2012, among others): a class of repairs whereby the speaker corrects themselves on the fly within the same conversational turn (e.g. “User: I want to go to London uhm sorry Paris”). Similarly, the crucial role of generating and responding to CRs (e.g. “Pardon/what/who?”) in conversational models has long been recognised (see San-Segundo et al., 2001; Purver, 2004; Purver and Ginzburg, 2004; Rieser and Moore, 2005; Rodríguez and Schlangen, 2004; Rieser and Lemon, 2006, among others), but existing systems either remain limited (e.g. Curry et al. (2018)) or do not support this at all – see Purver et al. (2018) for an overview. Recent work tries to identify when to pose a CR Addlesee et al. (2024); Madureira and Schlangen (2024, 2023); Zhu et al. (2021); Shi et al. (2022); Addlesee and Eshghi (2021), but few evaluate the ability of models to process their responses Gervits et al. (2021); Aliannejadi et al. (2021). Chiyah-Garcia et al. (2023) show that the ability of Vision and Language Models (VLM) to effectively process clarificational exchanges depends on the level of granularity of the model’s cross-modal representation, and therefore also the specific objectives used to train the models.
@@ 3 The BlockWorld-Repair Dataset
Prior works have struggled to handle the most ambiguous setting with blank blocks Tan and Bansal (2018); Mehta and Goldwasser (2019); Dan et al. (2021). However, here we argue that single-turn instructions are insufficient to solve highly ambiguous environments or real-world situations. Even the most detailed referring expressions in the Block World may fail to uniquely identify a referent (see Figure 1), which repairs could alleviate by narrowing down the candidate pool and introducing new information to identify the correct referent (T3 in Figure 1). Similar to how humans misinterpret and subsequently repair referential ambiguities with TPRs, VLMs must also be capable of handling multi-modal TPRs in dialogues, as they are essential for mutual coordination.
Differences from the original Block World
The original dataset consists of pairs of images and single-turn instructions, requiring agents to resolve inherently ambiguous commands. However, it does not account for dialogue or clarificational exchanges as strategies to overcome environmental ambiguity, which humans commonly use to repair miscommunications. To address this, we extend the dataset by incorporating brief dialogues where humans interact with the agent and produce Third Position Repairs. We preserve the original Block World instructions and test set while adding our dataset as an extension to explore how humans resolve ambiguity through clarification and repair strategies.
@@
In contrast to more recent photo-realistic simulated environments (e.g., AI2Thor Kolve et al. (2017)), we build BW-R using the Unity engine to simulate a realistic tabletop manipulation task that gives us the ability to specifically control for the conditions that require collaboration between agents to correctly complete the task. This is also similar to previous works that sacrifice photo-realistic vision to asses the systematic generalisation ability of VLMs for tabletop manipulation tasks (e.g., VIMA-Bench Jiang et al. (2023)), or works that focus on dialogue coordination in highly ambiguous scenarios (e.g., The Cups dataset Dobnik et al. (2020)). Appendix A provides further comparisons to related datasets.
@@ 3.1 Dataset Collection
As a result of this data collection, we obtained 795 dialogues after filtering out low-quality and problematic ones, which include 795 initial instructions, 629 source block TPRs and 635 final target position TPRs, resulting in 2059 total entries. We provide further details about the AMT setup, task and dataset statistics in Section B.1.
@@ 4.1.1 Dataset
We combine the Block World and BW-Rdata (70/30 train/test), which contains two types of entries: single-instructions (1 turn) and TPRs (3 turns: instruction, candidate response, repair). We ablate the data during our experiments to analyse fine-grained information about the models’ capabilities. Refer to Appendix B for further dataset details.
@@ 4.1.2 Tasks
Previous works have formulated these tasks to aid the prediction, such as finding a reference block and an offset distance from it Bisk et al. (2016) or quadrants in a board Mehta and Goldwasser (2019). For our experiments, and following common visual grounding objectives in VLMs Wang et al. (2022); Liu et al. (2024); Chen et al. (2023), we prompt models to generate the coordinates of the bounding box for the block or location ($[x_{min},y_{min},x_{max},y_{max}]$) in the range of 0-1, which we then scale to the image size. Predictions not matching this format during parsing are marked as out of distribution and removed. We use specific prompts for each model to match their training prompting strategy (see Appendix C).
@@ Results
Overall, the models struggle to exploit TPRs, especially in zero-shot settings. Intuitively, repairs should facilitate the task by introducing new information that refines the candidate pool. However, because repairs are often context-dependent (i.e., based on the candidate response), models require strong object-centric representations Bengio et al. (2013); Seitzer et al. (2023) and a deeper understanding of relational dynamics between objects to effectively leverage TPRs (e.g., “the block to the right”) Ilinykh and Dobnik (2021). This is a particular issue of current vision transformers, which frequently struggle with positional concepts such as spatial understanding Ilinykh and Dobnik (2022); Pantazopoulos et al. (2024), and tend to rely on language biases Salin et al. (2022). GPT-4o is unique in its ability to process these repairs effectively and improve its performance with instructions. We see that models can learn to process TPRs with some training, and the following section explores methods to further distil this capability.
@@ 5.2 Results
Table 2 shows the fine-tuned VLMs on the full data across different masking criteria [^4] All model results are provided in Appendix E.. We first observe that applying a mask on all tokens but completions has large benefits for both VLMs for repairs in source prediction, with strong improvements particularly when fine-tuning with repairs-only or the full dataset. This loss allows models to learn from TPRs and transfer some task knowledge from instructions to repairs when using the full data. In other settings, this loss has small negative effects on the VLM’s performance, especially when models are trained and tested on the same data. This is somewhat expected as calculating the loss for the complete input has benefits, such as using many more tokens during training, learning the input format and quickly adapting to the task. However, these models do not seem to learn the task well and thus do not generalise.
@@
Furthermore, we see that models fine-tuned on instructions alone, with or without masking, acquire some task capabilities, but do not generalise well to repairs, and vice-versa. When trained with the full data and completion-only loss, VLMs learn what is important regardless of the conversation, and reach their best results, learning what is important particularly for repairs.
@@ 6.1 Human Comparison
To compare models with humans on this task, we use the data samples that we collected from our in-person human study, in which participants saw the same data that models are evaluated on. The results in Table 3 show clear differences, with models struggling to reach human performance but demonstrating significant benefits from TPRs after training. We can see that, despite Idefics2 and LLaVA having lower source accuracy for instructions, they both beat GPT-4o on distance for source. In this case, accuracy fails as a metric for prediction quality, since lower distances better reflect true proximity. We also see impressive results for LLaVA processing TPRs after our training for source prediction. Regarding target position, this task is considerably more challenging for humans and models, reflected in worse scores overall. LLaVA and Idefics2 struggle to process repairs related to target locations, in contrast to GPT-4o which outperforms the human participants.
@@ 6.2 Impact of Task Difficulty
To better pinpoint the strengths and weaknesses of the different models, we categorise the BW-R data into difficulty levels according to human performance [^5] We provide this analysis according to GPT-4o performance in Appendix D.. We use lower human performance as an indirect proxy for the complexity of the example, and define three levels as follows: 1) Easy: both human annotators for the same entry correctly predicted the block (100% source accuracy) or their target position was within 1 block unit distance away from the true location; 2) Medium: at least one out of the two annotators correctly predicted the block (50% source accuracy) or their target position was below the human’s mean distance away from the location (between 1 and 3.22 units away); 3) Hard: neither of the annotators correctly predicted the block (0% source accuracy) or their target position was further than the human’s mean distance (above 3.22 units).
@@
The results of evaluating VLMs on these difficulty subsets (Table 4) indicate that models and humans do not struggle on the same entries. Models show a more uniform distribution of performance, with improvements over humans in the most difficult subset but significantly underperforming with easier references. This behaviour seems to follow the trend with the average word number in the subsets [^6] See Appendix D., suggesting that easy examples (most words) may employ longer, more complex terms to describe references, whereas hard examples (fewest words) may have easier terms for models, but are too underspecified for humans.
@@ 6.3 Qualitative Analysis
We further analyse 50 BW-R dialogues to understand the proficiency differences of VLMs with that of human participants (see Figure 3 for some examples). We find that humans and GPT-4o tend to have similar predictions when processing candidate responses and their repairs. Models are usually able to process simple spatial repairs (i.e., left, right, above) Chiyah-Garcia et al. (2023), however, they particularly struggle with more abstract concepts (i.e., rows, columns) Ilinykh et al. (2022). In that regard, both Idefics2 and LLaVA exhibit comprehension of the simpler referring expressions, but have a challenging time with longer or more complex sequences. Idefics2 also seems prone to over-correct (e.g., right pushes predictions to the absolute right). We also notice that VLMs sometimes predict bounding boxes that do not align with blocks in the source prediction task, which is unexpected given their pre-training (and fine-tuning) with referring expression objectives. For the target prediction task, VLMs significantly struggle and do not process repairs. GPT-4o is the only exception, which shows a more fine-grained understanding of the task yet falls behind on more difficult dialogues.
@@ 7 Conclusion
This paper explores VLM’s capabilities of handling repairs with a new dataset, BlockWorld-Repairs. We collect 795 collaborative dialogues in a highly ambiguous tabletop manipulation environment that requires complex multi-modal task instructions as well as repairs. In these scenarios, referential ambiguities are common and thus being able to process TPRs becomes essential to complete the task. We validate these dialogues through an in-person human study where participants attempt to solve the task, providing a baseline for model performance.
@@
We finish with an error analysis comparing these models and human participants, finding stark differences in how they process repairs. Surprisingly, models struggle with the dialogues that human participants found easiest, highlighting a large gap in their ability to process TPRs, despite these being fundamental to effective dialogue coordination. We believe that future research should focus on two major improvements for VLMs: 1) design training objectives that facilitate learning from interaction, including TPRs; 2) equip models with more object-centric visual representations that facilitate visual grounding tasks Parekh et al. (2024) which are of paramount importance for situated collaborative tasks.
@@ 8 Limitations
Another limitation of our work is that we do not attempt to model the entire spectrum of a robotic manipulation task (e.g., Octo Model Team et al., 2024). Instead, we follow previous work in Embodied AI (e.g., ALFRED Shridhar et al. (2020), Simbot Arena Gao et al. (2023), inter alia) which casts object manipulation as API call generation of an action label combined with a bounding box around the object to be manipulated. Although we don’t generate actions, the prediction tasks that we explore in BW-R involve generating the coordinates of a bounding box that can be used for picking up a specific block or putting it down onto a specific box on the board.